OpenStudy (waleed_imtiaz):

if the 1st term of an infinite geometric series is equal to twice of the sum of all the terms that follows it, then the value of r will be : ? A)1/3 B)1 C)1/2 D)2 please explain me the answer........!

OpenStudy (anonymous):

\[a _{1}=2S _{ \infty }\] \[S _{ \infty }=a/1-r\] solve

OpenStudy (waleed_imtiaz):

Can u tell me what U got after solving,.... ?

OpenStudy (anonymous):

r=1

OpenStudy (waleed_imtiaz):

Where's the mistake ?

OpenStudy (anonymous):

-1<r<1

OpenStudy (waleed_imtiaz):

okay.... BUt the answer on the book is 1/3... I dont know how it comes........... :(

OpenStudy (anonymous):

your working is correct, but the answer cant be -1or 1according to the restriction the problem is our equations the question says terms that 'follows it'

OpenStudy (waleed_imtiaz):

But @Jonask I here think of something...... U know the sum of infintite geometric series would be applicable only if |r|<1.... but if U got |r|>1. then the formula isn't applicable..... and the question I percept isn't really clear to me... please help me

OpenStudy (anonymous):

a and a1 are not the same, the term a will be \[a=a _{1}r\]

OpenStudy (anonymous):

its \[a _{1},a _{1} r,a _{1}r^2...\] \[S \infty=a/(1-r)=a _{1}r/1-r\]

OpenStudy (anonymous):

the terms following it are \[a _{1}r,a _{2}r^2,...\]

OpenStudy (waleed_imtiaz):

@Jonask the first term wouldn't be a1r.... Because U know the formula.. \[a_{n}=a_1r ^{n-1}\] so If U will put a=1.... then see what U will obtain.......

OpenStudy (anonymous):

these are the terms following (after) a

OpenStudy (waleed_imtiaz):

so I think the series should be like this...... \[a_1,a_1r ^{1},a_1r ^{2}......\]

OpenStudy (anonymous):

the series is like that but what about the series for terms after a1

OpenStudy (waleed_imtiaz):

after a...... every number is multiplied by r.... ?

OpenStudy (anonymous):

yes \[a _{1}=2S \infty\] \[S \infty =a/1-r\] and \[a=a _{1}r\]

OpenStudy (anonymous):

\[a _{1}=2[a/1-r]\] \[a _{1}=2a _{1}r/1-r\]

OpenStudy (anonymous):

\[1-r=2r\] \[3r=1\] \[r=1/3\]

OpenStudy (waleed_imtiaz):

So @Jonask in the formula..... what is 'a'? Its the second term or first......?

OpenStudy (waleed_imtiaz):

@Jonask ........ ?